# Geometric and topological rigidity of pinched submanifolds II

### v1

@@ 3 Proofs of the main results
Assume first that $\tilde{M}^{4}$ is diffeomorphic to $\mathbb{S}^{4}$. Arguing as in the proof of Case I, we conclude that $M^{4}$ is locally irreducible. Then, by [3, Th. 2], we have that $M^{4}$ is either diffeomorphic to a spherical space form or isometric to a compact symmetric space. In the former case, it is clear that the manifold $M^{4}$ must be diffeomorphic to $\mathbb{S}^{4}$. In the latter case, since the only 4-dimensional compact symmetric spaces are the round spheres, the product of two 2-dimensional spheres, and the complex or the quaternionic projective space $\mathbb{H}P^{1}$, it follows that $M^{4}$ must also be diffeomorphic to $\mathbb{S}^{4}$.
@@
Case III. Finally, suppose that the fundamental group $\pi_{1}(M)$ of $M^{4}$ is infinite. We first claim that $M^{4}$ is locally reducible. If this were not the case, then by Theorem 2 in [3], it would follow that the universal cover $\tilde{M}^{4}$ is diffeomorphic to a spherical space form, biholomorphic to $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{2}$, or isometric to a compact symmetric space. Each of these possibilities contradicts the fact that $\tilde{M}^{4}$ is not compact. Therefore, $M^{4}$ must be locally reducible. From this, it follows by Theorem 3.1 in [15] that the universal cover $\tilde{M}^{4}$ is isometric to one of the following:
@@
    $(\mathbb{R}^{k},g_{0})\times(N_{1}^{n_{1}},g_{1})\times(N_{2}^{n_{2}},g_{2})$, where $k\geq 0,g_{0}$ is the flat Euclidean metric, and either $n_{i}=2$ and $N_{i}=\mathbb{S}^{2}$ has nonnegative Gaussian curvature, or else $n_{i}=3$ and $N_{i}$ is compact with nonnegative Ricci curvature.
@@
First, we claim that case (ii) above cannot occur. Suppose, to the contrary, that it does. Clearly the immersion $\tilde{f}=f\circ\pi$ satisfies ($\ast$1), where $\pi\colon\tilde{M}^{4}\to M^{4}$ is the covering map. We consider the isometric immersions $f_{\varSigma}=\tilde{f}\circ i_{\varSigma}$ and $f_{N}=\tilde{f}\circ i_{N}$, where $i_{\varSigma}\colon\varSigma\to\tilde{M}$ and $i_{N}\colon N\to\tilde{M}$ are the inclusion maps. Since both immersions $i_{\varSigma},i_{N}$ are totally geodesic, it follows that the second fundamental form of $f_{\varSigma}$ and $f_{N}$ are given, respectively, by
$\alpha_{f_{\varSigma}}(X,Y)$ $=\alpha_{\tilde{f}}(i_{\varSigma_{*}}X,i_{\varSigma_{*}}Y),\;X,Y\in T\varSigma,$
$\alpha_{f_{N}}(V,W)$ $=\alpha_{\tilde{f}}(i_{N_{*}}V,i_{N_{*}}W),\;V,W\in TN.$
@@
$S_{\tilde{f}}$ $=S_{f_{\varSigma}}+S_{f_{N}}+\sum_{i=1,2}\sum_{j=3,4}\|\alpha_{\tilde{f}}(i_{\varSigma_{*}}e_{i},i_{N_{*}}e_{j})\|^{2},$ (18)
@@
$\|\Phi_{f_{\varSigma}}\|^{2}=S_{f_{\varSigma}}-4H^{2}_{f_{\varSigma}}\;{\text{and}}\;\|\Phi_{f_{N}}\|^{2}=S_{f_{N}}-4H^{2}_{f_{N}},$
respectively. Taking into account (18), (19) and the above, the inequality ($\ast$1) for $\tilde{f}$ can be equivalently written as
$\|\Phi_{f_{\varSigma}}\|^{2}+\|\Phi_{f_{N}}\|^{2}+2\sum_{i=1,2}\sum_{j=3,4}\|\alpha_{\tilde{f}}(i_{\varSigma_{*}}e_{i},i_{N_{*}}e_{j})\|^{2}\leq 4\big{(}c+{\langle}\mathcal{H}_{f_{\varSigma}},\mathcal{H}_{f_{N}}{\rangle}\big{)}.$ (20)
On the other hand, from the Gauss equation and ($\ast$1), it follows that the scalar curvature $\tau$ of $\tilde{M}^{4}$ satisfies $\tau\geq 4c+8\tilde{H}^{2}$. Since $\tau=2(K_{\varSigma}+K_{N})$, we then obtain the inequality
@@
Thus, only case (i) holds. Clearly $k\neq 3$. We now claim that $k=1$. If $k=4$, then $M^{4}$ is flat, and by the Gauss equation, we have $S=12c+16H^{2}$. This, together with ($\ast$1), implies that $c=0$ and $f$ is minimal, leading to a contradiction. If $k=2$, then the curvature operator of $M^{4}$ is nonnegative. Then it follows from Theorem 1.3 in [8] that $M^{4}$ is diffeomorphic to a quotient of one of the spaces $\mathbb{S}^{4}$, $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{4}$, $\mathbb{R}^{1}\times\mathbb{S}^{3}$, $\mathbb{S}^{2}\times\mathbb{S}^{2}$, $\mathbb{R}^{2}\times\mathbb{S}^{2}$, or $\mathbb{R}^{4}$ by a finite group of fixed-point-free isometries in the standard metric. This contradicts the fact that $\pi_{1}(M)$ is infinite. Therefore, $k=1$ and $\tilde{M}$ splits isometrically as ×$\mathbb{R}\times N^{3}$, where $N^{3}$ is compact, simply connected, and it has nonnegative Ricci curvature. From Theorem 1.2 in [8], we then conclude that $N^{3}$ is diffeomorphic to $\mathbb{S}^{3}$.∎
Proof of Corollary 3: Let $n\geq 5$, and suppose that the submanifold $f$ is not as described in parts $(i)$-$(iii)$ of Theorem 1. Then, the homology groups of $M^{n}$ satisfy
$H_{p}(M^{n};\mathbb{Z})=0\;\,\text{for}\;\,2\leq p\leq n-2,\;\,\text{and}\;\,H_{1}(M^{n};\mathbb{Z})\cong\mathbb{Z}^{\beta_{1}(M)}.$
@@
For the case $n=4$, the result follows directly from Theorem 2.∎
@@ 3.1 Examples of submanifolds satisfying condition ($\ast$1)
We now present a method for constructing geometrically distinct submanifolds that are diffeomorphic to either the sphere $\mathbb{S}^{n}$ or the torus $\mathbb{S}^{1}\times\mathbb{S}^{n-1}$, while also satisfying ($\ast$1) for $c=0$.
@@
Let $g\colon N^{n-1}\to\mathbb{R}^{m_{1}},n\geq 4$, be an isometric immersion of a manifold $N^{n-1}$ satisfying condition (italic-$\ast$1) for $k=\ell-1$, with $22\leq\ell\leq n-$ 2, as strict inequality at any point, namely,
@@
Consider any closed unit-speed curve $\gamma\colon\mathbb{S}^{1}\to\mathbb{R}^{m_{2}}$, whose first curvature $\kappa_{1}$ satisfies the inequality
@@
Then, the product immersion $f=\gamma\times g\colon\mathbb{S}^{1}\times N^{n-1}\to\mathbb{R}^{m_{1}+m_{2}}$ satisfies condition (italic-$\ast$1) for $k=\ell$.
@@
From these expressions, it follows that condition ($\ast$1) for $f$, with $k=\ell$ and $c=0$, is equivalent to the inequality (23), thereby completing the proof.∎
Next, we prove that there exist many isometric immersions $g\colon N^{n-1}\to\mathbb{R}^{m}$ that satisfy the conditions required in Proposition 11. Specifically, we provide geometrically distinct immersions of $\mathbb{S}^{n}$ into the Euclidean space $\mathbb{R}^{n+1}$ that satisfy ($\ast$1) for any $n\geq 3$ and $11\leq k\leq n-$ 1, either as a strict inequality or otherwise.
@@
for some $11\leq k\leq n-$ 1, then $f$ satisfies condition (italic-$\ast$1). If the inequality (24) is strict, then (italic-$\ast$1) is also strict.
@@
Since $a(n,k,H,0)=n^{2}H^{2}/(n-k)$, the proof follows directly from the inequalities $S\leq n\lambda^{2}_{n}$ and $H\geq\lambda_{1}$.∎
A large class of ellipsoids satisfies (24). Consider, for instance, the ellipsoid in $\mathbb{R}^{n+1}$ defined by
@@
where $0<a_{1}\leq\dots\leq a_{n+1}$. A straightforward computation shows that the minimum and the maximum of the principal curvatures of the ellipsoid are $a_{1}/a_{n+1}^{2}$ and $a_{n+1}/a_{1}^{2}$, respectively. It follows that condition (24) is satisfied if
@@
By the classical Hadamard theorem, any ovaloid in $\mathbb{R}^{n+1}$ is diffeomorphic to $\mathbb{S}^{n}$. Consequently, from Propositions 11 and 12, we derive geometrically distinct isometric immersions of manifolds that are diffeomorphic to either the sphere $\mathbb{S}^{n}$ or the torus $\mathbb{S}^{1}\times\mathbb{S}^{n-1},n\geq 4$, which also satisfy ($\ast$1) for $c=0$ and $k=2$. This establishes the optimality of our results.
@@
Consider a rotational hypersurface in $\mathbb{R}^{n+1},n\geq 4$, obtained by rotating a curve $c$ in the $x_{1}x_{2}$-plane around the $x_{1}$-axis. Suppose the curve is given as a graph of a positive function $u=u(x_{1})$. By adjusting the orientation, the rotational hypersurface has two principal curvatures given by
@@
with $\lambda$ having multiplicity $n-1$. The hypersurface satisfies the condition ($\ast$1) for $k=2$ as strict inequality if
@@
$-a_{n}(1+(u^{\prime})^{2})<uu^{\prime\prime}<b_{n}(1+(u^{\prime})^{2}),$
@@
By connecting such hypersurfaces to spheres as in [13, Example 3.3], we obtain compact hypersurfaces that satisfy the condition ($\ast$1) for $k=2$, with the first Betti number being any nonnegative integer.

### v2

@@ 3 Proofs of the main results
Assume first that $\tilde{M}^{4}$ is diffeomorphic to $\mathbb{S}^{4}$. Arguing as in the proof of Case I, we conclude that $M^{4}$ is locally irreducible. Then, by Theorem 2 in [3], $M^{4}$ is either diffeomorphic to a spherical space form or isometric to a compact symmetric space. In the former case, it is clear that the manifold $M^{4}$ must be diffeomorphic to $\mathbb{S}^{4}$. In the latter case, since the only 4-dimensional compact symmetric spaces are the round spheres, the product of two 2-dimensional spheres, the complex plane, and the quaternionic projective line $\mathbb{H}P^{1}$, it follows that $M^{4}$ must also be diffeomorphic to $\mathbb{S}^{4}$.
@@
Case III. Finally, suppose that the fundamental group $\pi_{1}(M)$ of $M^{4}$ is infinite. We first claim that $M^{4}$ is locally reducible. If this were not the case, then by Theorem 2 in [3], it would follow that either $M^{4}$ is diffeomorphic to a spherical space form, or its universal cover $\tilde{M}^{4}$ is biholomorphic to $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{2}$, or isometric to a compact symmetric space. Each of these possibilities contradicts the fact that $\tilde{M}^{4}$ is not compact. Therefore, $M^{4}$ must be locally reducible. It then follows from Theorem 3.1 in [18] that the universal cover $\tilde{M}^{4}$ is isometric to one of the following Riemannian products:
@@
    $(\mathbb{R}^{n_{0}},g_{0})\times(N_{1}^{n_{1}},g_{1})\times(N_{2}^{n_{2}},g_{2})$, where $n_{0}\geq 0,g_{0}$ is the flat Euclidean metric, and either $n_{i}=2$ and $N_{i}=\mathbb{S}^{2},i=1,2$, has nonnegative Gaussian curvature, or else $n_{i}=3$ and $N_{i}$ is compact with nonnegative Ricci curvature.
@@
First, we claim that case (ii) above cannot occur. Suppose, for the sake of contradiction, that it does. Clearly the immersion $\tilde{f}=f\circ\pi$ satisfies ($\ast$1), where $\pi\colon\tilde{M}^{4}\to M^{4}$ is the covering map. We consider the isometric immersions
$f_{\varSigma}=\tilde{f}\circ\mathfrak{i}_{\varSigma}\;\;\text{and}\;\;f_{N}=\tilde{f}\circ\mathfrak{i}_{N},$
where $\mathfrak{i}_{\varSigma}\colon\varSigma\to\tilde{M}$ and $\mathfrak{i}_{N}\colon N\to\tilde{M}$ are the inclusion maps. Since both immersions $\mathfrak{i}_{\varSigma},\mathfrak{i}_{N}$ are totally geodesic, it follows that the second fundamental form of $f_{\varSigma}$ and $f_{N}$ are given, respectively, by
$\alpha_{f_{\varSigma}}(X,Y)$ $=\alpha_{\tilde{f}}(\mathfrak{i}_{\varSigma_{*}}X,\mathfrak{i}_{\varSigma_{*}}Y),\;X,Y\in T\varSigma,$
$\alpha_{f_{N}}(V,W)$ $=\alpha_{\tilde{f}}(\mathfrak{i}_{N_{*}}V,\mathfrak{i}_{N_{*}}W),\;V,W\in TN.$
@@
$S_{\tilde{f}}$ $=S_{f_{\varSigma}}+S_{f_{N}}+2\sum_{i=1,2}\sum_{j=3,4}\|\alpha_{\tilde{f}}(\mathfrak{i}_{\varSigma_{*}}e_{i},\mathfrak{i}_{N_{*}}e_{j})\|^{2}$ (23)
and
@@
$\|\Phi_{f_{\varSigma}}\|^{2}=S_{f_{\varSigma}}-2H^{2}_{f_{\varSigma}}\;\;{\text{and}}\;\;\|\Phi_{f_{N}}\|^{2}=S_{f_{N}}-2H^{2}_{f_{N}},$
respectively. Taking into account (23), (24) and the above, inequality ($\ast$1) for $\tilde{f}$ can be equivalently written as
$\|\Phi_{f_{\varSigma}}\|^{2}+\|\Phi_{f_{N}}\|^{2}+2\sum_{i=1,2}\sum_{j=3,4}\|\alpha_{\tilde{f}}(\mathfrak{i}_{\varSigma_{*}}e_{i},\mathfrak{i}_{N_{*}}e_{j})\|^{2}\leq 4\big{(}c+{\langle}\mathcal{H}_{f_{\varSigma}},\mathcal{H}_{f_{N}}{\rangle}\big{)}.$ (25)
On the other hand, from the Gauss equation and ($\ast$1), it follows that the scalar curvature $\tau$ of $\tilde{M}^{4}$ satisfies
$\tau\geq 8(c+\tilde{H}^{2}).$
Since $\tau=2(K_{\varSigma}+K_{N})$, we then obtain the inequality
@@
Thus, only case (i) holds. Clearly, $n_{0}\neq 3$. We now claim that $n_{0}=1$. If $n_{0}=4$, then $M^{4}$ is flat. By the Gauss equation, we obtain
$S=12c+16H^{2}.$
Together with ($\ast$1), this implies that $c=0$ and $f$ is minimal, which leads to a contradiction. If $n_{0}=0$ or $n_{0}=2$, then the curvature operator of $M^{4}$ is nonnegative. In this case, by Theorem 1.3 in [10], $M^{4}$ is diffeomorphic to a quotient of one of the spaces $\mathbb{S}^{4}$, $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{4}$, $\mathbb{R}^{1}\times\mathbb{S}^{3}$, $\mathbb{S}^{2}\times\mathbb{S}^{2}$, $\mathbb{R}^{2}\times\mathbb{S}^{2}$, or $\mathbb{R}^{4}$ by a finite group of fixed-point-free isometries in the standard metric. This contradicts the fact that $\pi_{1}(M)$ is infinite. Therefore, $n_{0}=1$ and $\tilde{M}$ splits isometrically as $\mathbb{R}\times N^{3}$, where $N^{3}$ is compact, simply connected, and has nonnegative Ricci curvature. From Theorem 1.2 in [10], we then conclude that $N^{3}$ is diffeomorphic to $\mathbb{S}^{3}$.∎
Proof of Theorem 3: Let $n\geq 5$, and suppose that the submanifold $f$ is not as described in parts (i)(i)-$(iii)$ of Theorem 1. Then, the homology groups of $M^{n}$ satisfy
$H_{p}(M^{n};\mathbb{Z})=0\;\,\text{for all}\;\,2\leq p\leq n-2\;\;\text{and}\;\,H_{1}(M^{n};\mathbb{Z})\cong\mathbb{Z}^{\beta_{1}(M)}.$
@@
In the case $n=4$, the result follows directly from Theorem 2.∎
@@ 3.1 Examples of submanifolds satisfying condition ($\ast$1)
We now describe a method for constructing geometrically distinct submanifolds that are diffeomorphic either to the sphere $\mathbb{S}^{n}$ or to the torus $\mathbb{S}^{1}\times\mathbb{S}^{n-1}$, and which, in addition, satisfy condition ($\ast$1) in the case $c=0$.
@@
Let $g\colon N^{n-1}\to\mathbb{R}^{m_{1}},n\geq 4$, be an isometric immersion of a manifold $N^{n-1}$ satisfying condition ($\ast$1) for an integer $k=\ell-1$, with $2\leq\ell\leq n-2$, as a strict inequality at every point; that is,
@@
Consider any closed unit-speed curve $\gamma\colon\mathbb{S}^{1}\to\mathbb{R}^{m_{2}}$ whose first curvature $\kappa_{1}$ satisfies
@@
Then, the product immersion
$f=\gamma\times g\colon\mathbb{S}^{1}\times N^{n-1}\to\mathbb{R}^{m_{1}+m_{2}}$
satisfies condition ($\ast$1) for $k=\ell$.
@@
From these expressions, it follows that condition ($\ast$1) for $f$, with $k=\ell$ and $c=0$, is equivalent to inequality (28), thereby completing the proof.∎
Next, we show that there exist many isometric immersions $g\colon N^{n-1}\to\mathbb{R}^{m}$ satisfying the hypotheses of Proposition 11. In particular, we construct geometrically distinct immersions of $\mathbb{S}^{n}$ into the Euclidean space $\mathbb{R}^{n+1}$ that satisfy condition ($\ast$1) for all $n\geq 3$ and $1\leq k\leq n-1$, either with strict inequality or with equality.
@@
for some integer $1\leq k\leq n-1$, then $f$ satisfies condition ($\ast$1) at every point. Moreover, if inequality (29) is strict, then inequality ($\ast$1) is also strict at every point.
@@
Since
$a(n,k,H,0)=\frac{n^{2}H^{2}}{n-k},$
the proof follows directly from the inequalities $S\leq n\lambda^{2}_{n}$ and $H\geq\lambda_{1}$.∎
A large class of ellipsoids satisfies condition (29). Consider, for instance, the ellipsoid in $\mathbb{R}^{n+1}$ defined by
@@
where $0<a_{1}\leq\dots\leq a_{n+1}$. A straightforward computation shows that the minimum and the maximum of the principal curvatures of the ellipsoid are $a_{1}/a_{n+1}^{2}$ and $a_{n+1}/a_{1}^{2}$, respectively. It follows that condition (29) holds provided
@@
By Hadamard’s classical theorem, every ovaloid in $\mathbb{R}^{n+1}$ is diffeomorphic to $\mathbb{S}^{n}$. Consequently, Propositions 11 and 12 provide geometrically distinct isometric immersions of manifolds diffeomorphic either to the sphere $\mathbb{S}^{n}$ or to the torus $\mathbb{S}^{1}\times\mathbb{S}^{n-1}$ for $n\geq 4$, which also satisfy condition ($\ast$1) with $c=0$ and $k=2$. Thus, we obtain numerous compact, geometrically distinct submanifolds that strictly satisfy ($\ast$1) at every point. Moreover, this strict form is preserved under sufficiently small smooth deformations of any such example. Hence our results are optimal.
@@
Consider a rotational hypersurface in $\mathbb{R}^{n+1},n\geq 4$, obtained by rotating a curve $c$ in the $x_{1}x_{2}$-plane around the $x_{1}$-axis. Suppose the curve is given as a graph of a positive function $u=u(x_{1})$. After choosing the orientation appropriately, the rotational hypersurface has two principal curvatures given by
@@
where $\lambda$ has multiplicity n1n-1. The hypersurface satisfies condition ($\ast$1) with $k=2$ in strict form provided that
@@
$a_{n}(1+(u^{\prime})^{2})<uu^{\prime\prime}<b_{n}(1+(u^{\prime})^{2}),$
@@
By connecting such hypersurfaces to spheres, as in Example 3.3 in [16], we obtain compact hypersurfaces that satisfy condition ($\ast$1) with $k=2$, whose first Betti number can be any nonnegative integer. The homology of these hypersurfaces is precisely as described in part (iii) of Theorem 3.